From pdbforce at jlab.org Sun Jun 1 10:00:02 2025 From: pdbforce at jlab.org (pdbforce at jlab.org) Date: Sun, 1 Jun 2025 10:00:02 -0400 (EDT) Subject: [Halla_running] [New Logentry] RC Update -- Sunday 06/01/25 [NO Meeting] Message-ID: <20250601140002.485A2803684@elog9.acc.jlab.org> Logentry Text: --

Accelerator and Hall Status

Links to Shift Summaries

Sources of Major Downtime in the Hall

Source In the Last 24 hrs Since Wednesday May 28 Swing
DAQ 2.2 hrs 20.2 hrs
ECAL AC 0 hr 1 hr

Accumulated Charge on LH2

Run Plan for Today

We will not meet today. Please call the RC phone if you need me. Otherwise, have a great weekend!

--- This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/4388177 -------------- next part -------------- An HTML attachment was scrubbed... URL: From pdbforce at jlab.org Mon Jun 2 10:30:02 2025 From: pdbforce at jlab.org (pdbforce at jlab.org) Date: Mon, 2 Jun 2025 10:30:02 -0400 (EDT) Subject: [Halla_running] [New Logentry] RC Update -- Monday 06/02/25 [Meeting at 1:00pm] Message-ID: <20250602143002.6064F803836@elog9.acc.jlab.org> Logentry Text: --

Accelerator and Hall Status

Links to Shift Summaries

Sources of Major Downtime in the Hall

Source In the Last 24 hrs Since Wednesday May 28 Swing
DAQ 3.2 hrs 23.4 hrs
ECAL AC 0.5 hr 1.5 hrs

Accumulated Charge on LH2

Run Plan for Today

Let's meet today at the usual time (1:00 pm) to go over recent and future activities.
Meeting details:
Time: Daily at 1:00pm
Location: Counting house conference room
Join ZoomGov Meeting
https://jlab-org.zoomgov.com/j/1616220064
Meeting ID: 161 622 0064
Passcode: 314159

--- This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/4388862 -------------- next part -------------- An HTML attachment was scrubbed... URL: From pdbforce at jlab.org Tue Jun 3 10:35:01 2025 From: pdbforce at jlab.org (pdbforce at jlab.org) Date: Tue, 3 Jun 2025 10:35:01 -0400 (EDT) Subject: [Halla_running] [New Logentry] RC Update -- Tuesday 06/03/25 [Meeting at 1:00pm] Message-ID: <20250603143501.A6FA080368E@elog9.acc.jlab.org> Logentry Text: --

Accelerator and Hall Status

Links to Shift Summaries

Sources of Major Downtime in the Hall

Source In the Last 24 hrs Since Wednesday May 28 Swing
DAQ 4 hrs 27.4 hrs
ECAL AC 0.5 hr 1.5 hrs

Accumulated Charge on LH2

Run Plan for Today

Let's meet today at the usual time (1:00 pm) to go over recent and future activities.
Meeting details:
Time: Daily at 1:00pm
Location: Counting house conference room
Join ZoomGov Meeting
https://jlab-org.zoomgov.com/j/1616220064
Meeting ID: 161 622 0064
Passcode: 314159

--- This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/4389621 -------------- next part -------------- An HTML attachment was scrubbed... URL: From pdbforce at jlab.org Wed Jun 4 11:35:01 2025 From: pdbforce at jlab.org (pdbforce at jlab.org) Date: Wed, 4 Jun 2025 11:35:01 -0400 (EDT) Subject: [Halla_running] [New Logentry] RC Update -- Wednesday 06/04/25 [Short Meeting at 1:00pm] Message-ID: <20250604153501.E1DEF803684@elog9.acc.jlab.org> Logentry Text: --

Accelerator and Hall Status

  • Beam continued to be stable.
  • Early day shift we made an effort to raise the beam current and settled at 22.5uA. Unfortunately, around 3:30pm we had to stop data taking temporarily as the CH disks were full! Alex could free up some space quickly allowing us to continue data taking at 18 uA. Then around 6pm the disks were full again! We reduced the beam current to 11uA to keep up. Then around 8pm we raised it to 15uA and stayed there overnight. Ole is working with the computer center to resolve this issue.
  • Today is maintenance day. Beam was taken away at 6:45 am. Hall went to controlled access around 8:20am. Currently, experts are working in the hall to troubleshoot pending issues. The beam is expected to be back at 3pm.

Links to Shift Summaries

Sources of Major Downtime in the Hall

Source In the Last 24 hrs Since Wednesday May 28 Swing
DAQ 3.5 hrs 30.6 hrs
ECAL AC 0.5 hr 1.5 hrs

Accumulated Charge on LH2

  • Total Accumulated Charge at Kin3 so far: 15.53 C (Goal: 100 C)
  • Total Accumulated Charge at Kin3 since Wednesday May 28 Swing: 7.90 C

Run Plan for Today

  • Restore beam in the hall
  • Take production data at @ 22.5 uA. In case the file transfer speed is still an issue then reduce the beam current to accommodate for it.

Let's meet today at the usual time (1:00 pm) to go over recent and future activities.
Meeting details:
Time: Daily at 1:00pm
Location: Counting house conference room
Join ZoomGov Meeting
https://jlab-org.zoomgov.com/j/1616220064
Meeting ID: 161 622 0064
Passcode: 314159

--- This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/4390394 -------------- next part -------------- An HTML attachment was scrubbed... URL: From puckett at jlab.org Thu Jun 5 12:58:15 2025 From: puckett at jlab.org (Andrew Puckett) Date: Thu, 5 Jun 2025 16:58:15 +0000 Subject: [Halla_running] RC meeting reminder Message-ID: Hello all, This is your friendly reminder that we will hold the daily RC meeting today on Zoom and in the counting house 2nd floor conference room. Zoom meeting info is below: Join ZoomGov Meeting https://jlab-org.zoomgov.com/j/1616220064 Meeting ID: 161 622 0064 Passcode: 314159 https://sbs.jlab.org/cgi-bin/DocDB/private/ShowDocument?docid=431 We?ll discuss the last 24 hours running and the plan for the next 24 hours. Best regards, Andrew Puckett Professor, Physics Department University of Connecticut 196 Auditorium Road, Unit 3046 Storrs, CT 06269-3046 Office phone: (860) 486-7137 https://puckett.physics.uconn.edu puckett at jlab.org -------------- next part -------------- An HTML attachment was scrubbed... URL: From puckett at jlab.org Fri Jun 6 13:01:06 2025 From: puckett at jlab.org (Andrew Puckett) Date: Fri, 6 Jun 2025 17:01:06 +0000 Subject: [Halla_running] GEP RC meeting reminder today, 1 PM Message-ID: Hello all, This is a reminder for our daily experiment meeting starting at the usual time of 1 PM. I?d send you the link to the website, but my laptop can?t access the wiki page right now. The link can be found at the SBS wiki page. Meeting is starting. Best regards, Andrew -------------- next part -------------- An HTML attachment was scrubbed... URL: From puckett at jlab.org Mon Jun 9 12:33:02 2025 From: puckett at jlab.org (Andrew Puckett) Date: Mon, 9 Jun 2025 16:33:02 +0000 Subject: [Halla_running] Hall A RC update and meeting reminder, June 9, 2025 Message-ID: Hello all, Hall A continued with relatively smooth production running at 15 uA over the weekend. Overall running efficiency was slightly below the recent norm but still above 50%. Major reasons for significant downtime were accelerator and/or DAQ-related. For more details, see the RC update and run plan for June 9 at the wiki page here: https://sbs.jlab.org/wiki/index.php/RC_meeting_and_Run_Plan,_Monday,_June_9th We will have our daily RC meeting on Zoom (link at the wiki page linked above) and in person in the conference room on the 2nd floor of the counting house, at the usual time of 1:00 PM. Best regards, Andrew Andrew Puckett Professor, Physics Department University of Connecticut 196 Auditorium Road, Unit 3046 Storrs, CT 06269-3046 Office phone: (860) 486-7137 https://puckett.physics.uconn.edu puckett at jlab.org -------------- next part -------------- An HTML attachment was scrubbed... URL: From puckett at jlab.org Tue Jun 10 12:52:34 2025 From: puckett at jlab.org (Andrew Puckett) Date: Tue, 10 Jun 2025 16:52:34 +0000 Subject: [Halla_running] RC daily update and meeting reminder, June 10, 2025 Message-ID: Hello all, Hall A running was relatively smooth over the last 24 hours (particularly during day and swing shifts). We are now up to 4.2 C collected since the start of my RC tenure on June 4 and 19.8 C in total. Two significant changes to running conditions occurred since yesterday?s meeting: 1. HCAL trigger threshold was lowered to 20% of elastic kinetic energy due to concerns over energy calibration/threshold uncertainty/trigger efficiency 2. ECAL trigger threshold was increase to 65% of elastic, which should have a minimal impact on the trigger efficiency while significantly reducing the rate. As the big full replays with tracking started on Sunday evening start to pour in, new concerns have developed over GEM HV settings/gain/tracking efficiency. We are temporarily running at the relatively calmer beam current of 7 uA until we can systematically determine what are the optimal HV settings at which to run the experiment. Let?s meet on Zoom and in person at the CH 2nd floor conference room to discuss the run plan and the plans for tomorrow?s accelerator down. For more details and Zoom link info, see the RC daily meeting page at the wiki here: https://sbs.jlab.org/wiki/index.php/RC_meeting_and_Run_Plan,_Tuesday,_June_10th Best regards, Andrew Andrew Puckett Professor, Physics Department University of Connecticut 196 Auditorium Road, Unit 3046 Storrs, CT 06269-3046 Office phone: (860) 486-7137 https://puckett.physics.uconn.edu puckett at jlab.org -------------- next part -------------- An HTML attachment was scrubbed... URL: From puckett at jlab.org Wed Jun 11 12:48:50 2025 From: puckett at jlab.org (Andrew Puckett) Date: Wed, 11 Jun 2025 16:48:50 +0000 Subject: [Halla_running] Hall A RC daily update and meeting reminder, June 11, 2025 Message-ID: Hello all, This is a reminder for our daily RC meeting on Zoom and in the 2nd floor CH conference room at the usual time of 1 PM. Over the last 24 hours, we completed the 7 uA runs for the GEM HV studies and then resumed production at 15 uA during owl. Unfortunately DAQ issues developed during owl that kept Hall A down from around 4:30 until MCC took the beam away for their planned downtime. Hall is in controlled access during the down, and DAQ experts are in the hall fixing the ROC23 issue. For more details and Zoom info, see here: https://sbs.jlab.org/wiki/index.php/RC_meeting_and_Run_Plan,_Wednesday,_June_11th Best regards, Andrew Andrew Puckett Professor, Physics Department University of Connecticut 196 Auditorium Road, Unit 3046 Storrs, CT 06269-3046 Office phone: (860) 486-7137 https://puckett.physics.uconn.edu puckett at jlab.org -------------- next part -------------- An HTML attachment was scrubbed... URL: From heinricn at jlab.org Mon Jun 16 13:35:02 2025 From: heinricn at jlab.org (heinricn at jlab.org) Date: Mon, 16 Jun 2025 13:35:02 -0400 (EDT) Subject: [Halla_running] [New Logentry] Planned Access for Tuesday Jun 17 Message-ID: <20250616173502.8D090803833@elog9.acc.jlab.org> Logentry Text: -- The plan is for hall A to be placed in controlled access tomorrow (Jun 17). I just got off the phone with RAD CON, they explained that Hall A is 4th in the order and there will be 3 sweep teams starting at 7:45. He could not give me a firm time estimate, but would let me know when the sweep starts. Rad Con will contact Lawrence before they go in to lock out the SBS. I expect that we will be in controlled for most of day shift. Currently planned work items are: - Don Jones wants to install some shielding, he'll have some techs with him (3 Keys) - GEM Crew would like to replace a MPC Card on GEM bunker (3 Keys) - A few Items need to be surveyed by rad con to be removed from the hall (subject to Rad Con Availability) - ECAL crew want to check some channels (2 Keys) Let the Shift lead know before you go in the hall, and let me know if there is any other work that needs doing. Thanks, Nathan Heinrich --- This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/4399121 -------------- next part -------------- An HTML attachment was scrubbed... URL: From heinricn at jlab.org Tue Jun 17 09:20:01 2025 From: heinricn at jlab.org (heinricn at jlab.org) Date: Tue, 17 Jun 2025 09:20:01 -0400 (EDT) Subject: [Halla_running] [New Logentry] RC Update June 17 Message-ID: <20250617132001.DDA37803BC3@elog9.acc.jlab.org> Logentry Text: -- A quick update. The hall is in controlled, everyone who had work to do should be getting to it. I'd like work to be planned to finish by 2pm so that we can unlock the SBS by 3pm. Please let the shift crew know before you go down. Last night we had network problems with both the GEM and ECAL HV supplies. The GEMs required an access to fix, done at ~9pm last night. It was decided that we would wait until today to fix the ECAL because there was only 12 bars affected. ECAL fixing is now added to the list of work in the hall today. Also, big announcement at the PD meeting this morning is that tomorrow Day shift (Wednesday the 18th) will be beam studies for Moller. This means they will be sending test beam into the hall. So no Production data will be available for most of day shift. Thanks, Nathan --- This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/4399731 -------------- next part -------------- An HTML attachment was scrubbed... URL: From heinricn at jlab.org Wed Jun 18 11:55:02 2025 From: heinricn at jlab.org (heinricn at jlab.org) Date: Wed, 18 Jun 2025 11:55:02 -0400 (EDT) Subject: [Halla_running] [New Logentry] RC Update June 18 Message-ID: <20250618155502.9C7BD803701@elog9.acc.jlab.org> Logentry Text: -- Beam studies today, should begin restore at 1pm with beam back at 2pm. The Hot Channel fix yesterday appears to have reduced the trigger rate in the ECAL from 2kHz to 1.4kHz. ECAL calibration is implemented the experts are wait for beam to return to check it. Join RC meeting at 1:00pm https://sbs.jlab.org/cgi-bin/DocDB/private/ShowDocument?docid=431 Thanks, Nathan Heinrich --- This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/4400417 -------------- next part -------------- An HTML attachment was scrubbed... URL: From covrig at jlab.org Thu Jun 19 11:44:44 2025 From: covrig at jlab.org (Silviu Covrig Dusa) Date: Thu, 19 Jun 2025 15:44:44 +0000 Subject: [Halla_running] RC update for 19-Jun-2025 Message-ID: Due to the holiday there will not be a daily 1pm RC meeting today. If you have any issues or concerns regarding the run plan, please call the RC phone. Key updates for the last 24 h: * Beam back from beam studies around 12:50 yesterday, cross check, beam raster and positions checked fine, went to production * Machine was down for about 2h, during Swing, due to 2 cavities dropping out of HV (2L18 and 1L06, 2L18 was bypassed) * 8h run-plan with HCAL threshold at 10% and ECAL at 70%, beam current 15 ?A completed Swing-Owl * GEM baseline studies during Swing, baseline adjusted from 400-500 ADC ch to 800-900 ADC ch, summary in https://logbook.jlab.org/entry/44090882 * Since run 4770, 6:30am today, we are running at 22 ?A beam current with HCAL at 20% and ECAL at 70% and new GEM baseline, the plan is to stay here until further notice * Total Accumulated Charge at Kin3 so far: 25.82 C (Goal: 100 C) * Total Accumulated Charge at Kin3 since June 18 Swing: 0.9 C Run plan: Production on 30 cm LH2 cell at 22 uA with the current settings for GEMs, ECAL and HCAL Shift summaries: 18-Jun-2025 D https://logbooks.jlab.org/entry/4400380 18-Jun-2025 S https://logbooks.jlab.org/entry/4400629 19-Jun-2025 O https://logbooks.jlab.org/entry/4400873 Shift BTA: 18-Jun-2025 D https://logbooks.jlab.org/entry/4400609 18-Jun-2025 S https://logbooks.jlab.org/entry/4400869 19-Jun-2025 O https://logbooks.jlab.org/entry/4401133 Kind Regards, Silviu RC phone 757-876-1787 -------------- next part -------------- An HTML attachment was scrubbed... URL: -------------- next part -------------- A non-text attachment was scrubbed... Name: GEp5_24hBeamPlot_1819Jun2025.jpg Type: image/jpeg Size: 261683 bytes Desc: GEp5_24hBeamPlot_1819Jun2025.jpg URL: From covrig at jlab.org Fri Jun 20 09:55:02 2025 From: covrig at jlab.org (covrig at jlab.org) Date: Fri, 20 Jun 2025 09:55:02 -0400 (EDT) Subject: [Halla_running] [New Logentry] RC Report for 20-Jun-2025 Message-ID: <20250620135502.EBF90803BE3@elog9.acc.jlab.org> Logentry Text: -- Key updates for the last 24h: - Beam delivery has been excellent, average trip rate 7/h over the past 24h (as reported at the 8am MCC meeting) - MCC to decide if they will attempt to repair 2L18 today, before the weekend. The issue: there is no head room for the SL. If any other cavity goes down in SL there may be significant down time for repairs. - Riad announced that the current laser spot is getting to the tail end of its lifetime and a spot move is planned over the next 7-10 days. Should we consider a Moller measurement before the spot move? - HCAL threshold changed to 15%, per Mark's recommendation, running at this setting since about 11am 19-Jun-25. - As for the previous 24h, during this period too most of the experimental down time was caused by DAQ crashes. If I count from BTAs we lost 6-7h of beam time over the past 24h due to DAQ! Can something be done about it? - Total Accumulated Charge at Kin3 so far: 26.8 C (Goal: 100 C) - Total Accumulated Charge at Kin3 since June 18 Swing: 1.8 C Run plan: Production on the 30 cm LH2 target cell (loop 1) at 22 ?A, with the current settings for GEMs, ECAL and HCAL (unless subsystem experts want to change them). Shift summaries: 19-Jun-2025 D https://logbooks.jlab.org/entry/4401145 19-Jun-2025 S https://logbooks.jlab.org/entry/4401359 20-Jun-2025 O https://logbooks.jlab.org/entry/4401558 Shift BTA: 19-Jun-2025 D https://logbooks.jlab.org/entry/4401507 19-Jun-2025 S https://logbooks.jlab.org/entry/4401551 20-Jun-2025 O https://logbooks.jlab.org/entry/4401818 RC Daily Meetings1:00 PM , via Zoom and in person in the counting house meeting room (2nd floor). No meeting planned for the weekend, unless something important comes up. Zoom info: Join ZoomGov Meeting https://jlab-org.zoomgov.com/j/1616220064 Meeting ID: 161 622 0064 Passcode: 314159 --- This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/4401888 -------------- next part -------------- An HTML attachment was scrubbed... URL: From covrig at jlab.org Sat Jun 21 13:30:02 2025 From: covrig at jlab.org (covrig at jlab.org) Date: Sat, 21 Jun 2025 13:30:02 -0400 (EDT) Subject: [Halla_running] [New Logentry] RC Report for 21-Jun-2025 Message-ID: <20250621173002.19F9D803BDC@elog9.acc.jlab.org> Logentry Text: -- Key updates for the last 24h: - CC and PD are happy, beam delivery has been good over the past 24h. Major beam downtime for 2.8h during Day on Fri to repair 2L18 today. Beam back for production since 14:25 on Fri. - On our side, we are still plagued by daq crashes and recoveries, of the 14.5 ABUs for the last 24h, we wasted 5.24h on daq crash/recovery or 36% of beam time! - HCAL threshold changes during the weekend to assess if we can bump the beam current higher than 22 ?A on Mon - Total Accumulated Charge at Kin3 so far: 27.99 C (Goal: 100 C) - Total Accumulated Charge at Kin3 since June 18 Swing: 2.99 C Run plan: Production on the 30 cm LH2 target cell (loop 1) at 22 ?A, also on the wiki: https://sbs.jlab.org/wiki/index.php/RC_meeting_and_Run_Plan,_Saturday,_June_21st Current issues: - Over a 24h period we seem to waste more than 30% of beam-on-target with daq crashes and recoveries, which is an enormous penalty on statistics. Yesterday, during the daily RC meeting, it was suggested by the daq expert to reboot all crates in between two 4h runs, believing that such a procedure would make daq more stable. A daq "panic button" (https://logbooks.jlab.org/entry/4402072) was implemented and SLs were instructed to use it at the end of each 4h run, before starting a new run. Two things have become apparent since Swing 20-Jun-2025: the panic button does not always work as intended (https://logbooks.jlab.org/entry/4402271), and the daq is not more stable during a 4h run, still crashing with various issues. Today's Owl shift wasted 42.5% of beam time on daq issues. - Gentle reminder for shift-takers: Please put the RC phone and HA CH phone numbers (SL x5501, TO x5507) on you phone and please call the RC if you cannot make it on time for your shift. During summer traffic can be heavy accessing the lab. Shift summaries: 20-Jun-2025 D https://logbooks.jlab.org/entry/4401825 20-Jun-2025 S https://logbooks.jlab.org/entry/4402157 21-Jun-2025 O https://logbooks.jlab.org/entry/4402320 Shift BTA: 20-Jun-2025 D https://logbooks.jlab.org/entry/4402117 20-Jun-2025 S https://logbooks.jlab.org/entry/4402293 21-Jun-2025 O https://logbooks.jlab.org/entry/4402554 If you have an issue pertinent to operations, please reach out on the RC phone: 757-876-1787. No RC meeting today and tomorrow, enjoy your weekend! --- This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/4402683 -------------- next part -------------- An HTML attachment was scrubbed... URL: From covrig at jlab.org Sun Jun 22 13:05:02 2025 From: covrig at jlab.org (covrig at jlab.org) Date: Sun, 22 Jun 2025 13:05:02 -0400 (EDT) Subject: [Halla_running] [New Logentry] RC Report for 22-Jun-2025 Message-ID: <20250622170502.8CE7E803BC3@elog9.acc.jlab.org> Logentry Text: -- Key updates for the past 24h ( Sat Day and Swing and Sun Owl): - CC and PD: quiet past 24h, beam delivery was 22-of-24h for physics. Major beam downtime for ~40min this morning around 7:30am with a cavity trip. - On our side, Day and Swing lost about 2h combined on daq crashes and recovery, while Owl lost 4h on daq crashes and recovery! Over the past 24h we burned 30% of beam-on-target time unable to take data due to daq crashes and recovery! - HCAL threshold changes during the weekend to assess if we can bump the beam current higher than 22 ?A on Mon - SL please note that the primary daq expert, Alexander, is on vacation until Jul 10. During this time primary daq on-call will be C.H. Leung (217-693-1360), and second line will be Chandan (757-342-4035). - Total Accumulated Charge at Kin3 so far: 29.18 C (Goal: 100 C) - Total Accumulated Charge at Kin3 since June 18 Swing: 4.18 C Run plan: Production on the 30 cm LH2 target cell (loop 1) at 22 ?A, also on the wiki: https://sbs.jlab.org/wiki/index.php/RC_meeting_and_Run_Plan,_Sunday,_June_22nd Current issues: - Plagued still by daq crashes, during the last 24h we seem to have wasted 30% of beam time on daq crashes and recovery. The 4h of beam time lost during Owl shift seem to have been caused by an unresponsive power strip in the Hall. Once the crates connected to it were commented out in software, we were able to take data for the rest of Owl. Ching Him (SL for Day today) took an opportunistic access to the Hall, power cycled the strip locally and the crates connected to it seem to be able to take data: https://logbooks.jlab.org/entry/4403280 - APV baseline sagging/broadening: HV test showed a significant reduction of the width at lower gain. - Online zero suppression performance (common-mode min values for online zero suppression were changed to zero for the first iteration of averaging until we are ready to deploy ROI readout and histogramming firmware). - HCAL calibration/trigger efficiency, almost complete, still waiting on replays to finish. Shift summaries: 21-Jun-2025 D https://logbooks.jlab.org/entry/4402562 21-Jun-2025 S https://logbooks.jlab.org/entry/4402766 22-Jun-2025 O https://logbooks.jlab.org/entry/4402956 (grim reading) Shift BTA: 21-Jun-2025 D https://logbooks.jlab.org/entry/4402759 21-Jun-2025 S https://logbooks.jlab.org/entry/4402954 22-Jun-2025 O https://logbooks.jlab.org/entry/4403215 If you have an issue pertinent to operations, please reach out on the RC phone: 757-876-1787. No Rc daily meeting today, enjoy your weekend! --- This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/4403352 -------------- next part -------------- An HTML attachment was scrubbed... URL: From covrig at jlab.org Mon Jun 23 11:25:02 2025 From: covrig at jlab.org (covrig at jlab.org) Date: Mon, 23 Jun 2025 11:25:02 -0400 (EDT) Subject: [Halla_running] [New Logentry] RC Report for 23-Jun-2025 Message-ID: <20250623152502.0530C803701@elog9.acc.jlab.org> Logentry Text: -- Key updates for the past 24h (Sun Day+Swing and Mon Owl): - 17.93h actual beam delivery on target. The beam was down for about 6h with major culprits RF and vacuum issues. - On our side we were down for less than 2h during beam delivery, which accounts for 9.5% of wasted beam time, a big improvement on previous recent 24h periods. - Opportunistic access by SL CH Leung during Day on Sun to power cycle a power strip, which improved coms in daq: https://logbooks.jlab.org/entry/4403280 - DAQ on-call until Jul 10: primary C. H. Leung (217-693-1360), secondary Chandan (757-342-4035). - Total Accumulated Charge at Kin3 so far: 30.49 C (Goal: 100 C) - Total Accumulated Charge at Kin3 since June 18 Swing: 5.49 C Run plan: Continue production on the 30 cm LH2 target cell (loop 1) at 22 ?A. Current issues: - The last 24h have been quieter with respect to daq crashes, we only have 9.5% of "deadtime" in this period. The beam was also off during Owl, 2-3am, this may have helped our numbers some. - APV baseline sagging/broadening: HV test showed a significant reduction of the width at lower gain. - Online zero suppression performance (common-mode min values for online zero suppression were changed to zero for the first iteration of averaging until we are ready to deploy ROI readout and histogramming firmware). - HCAL calibration/trigger efficiency, almost complete, still waiting on replays to finish. Shift summaries: 22-Jun-2025 D https://logbooks.jlab.org/entry/4403223 22-Jun-2025 S https://logbooks.jlab.org/entry/4403448 23-Jun-2025 O https://logbooks.jlab.org/entry/4403639 Shift BTA: 22-Jun-2025 D https://logbooks.jlab.org/entry/4403426 22-Jun-2025 S https://logbooks.jlab.org/entry/4403633 23-Jun-2025 O https://logbooks.jlab.org/entry/4403867 We will have our daily RC meeting at 1pm in the conference room on the second floor of the CH and on zoom. Zoom info: Join ZoomGov Meeting https://jlab-org.zoomgov.com/j/1616220064 Meeting ID: 161 622 0064 Passcode: 314159 --- This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/4403979 -------------- next part -------------- An HTML attachment was scrubbed... URL: From covrig at jlab.org Tue Jun 24 11:30:02 2025 From: covrig at jlab.org (covrig at jlab.org) Date: Tue, 24 Jun 2025 11:30:02 -0400 (EDT) Subject: [Halla_running] [New Logentry] RC Report for 24-Jun-2025 Message-ID: <20250624153002.9781B803701@elog9.acc.jlab.org> Logentry Text: -- Key updates for the past 24h (Mon Day+Swing and Tue Owl): - 16.2h actual beam delivery on target. The beam was down for almost 8h with major culprits NL dropping to restricted (https://logbooks.jlab.org/entry/4403924) and cooling water chillers (https://logbooks.jlab.org/entry/4404500) - On our side we were down for 3.6h during beam delivery, which accounts for 22.2% of wasted beam time. Most of the downtime happened, again, during Owl, 2.2h, and during the first half of Owl, when beam was available. - Tomorrow Day will be beam studies 7-9am, then RF recovery until mid afternoon with beam recovery after. Access to the Hall can be scheduled after 9am tomorrow. We plan to go to CA. - Thu Day there are plans for beam studies during beam delivery to test two settings of the Wien filter (for MOLLER). The Halls are "required" to take data during this time, maybe tag it, and report about the effects (if any). MCC will announce when the Wien tests will start and stop, to give Halls time to prepare for daq. - DAQ on-call until Jul 10: primary C. H. Leung (217-693-1360), secondary Chandan (757-342-4035). - Total Accumulated Charge at Kin3 so far: 31.61 C (Goal: 100 C) - Total charge delivered on target since Jun 18 noon: 9.31 C - Total Accumulated Charge from the run list at Kin3 since June 18 noon: 6.61 C - Charge acquisition efficiency 71% (from the plot attached) Run plan: Continue production on the 30 cm LH2 target cell (loop 1) at 22 ?A and also on wiki: https://sbs.jlab.org/wiki/index.php/RC_meeting_and_Run_Plan,_Tuesday,_June_24th#Current_Issues Current Issues: - DAQ crashes: the last 24h we wasted 3.6h due to them or 22.2% of actual beam delivery. - Online zero suppression performance (common-mode min values for online zero suppression were changed to zero for the first iteration of averaging until we are ready to deploy ROI readout and histogramming firmware). - HCAL calibration/trigger efficiency, almost complete, still waiting on replays to finish. - Andrew pointed out that the sbb magnet current recorded in the run list changed on Jun 13 from 1762 A to 1757 A. There are two epics for the magnet's current, the setpoint (MSUPERBIGBITE) and the readback (MSUPERBIGBITEM). The current in the magnet, according to Ellen, is the readback. Both epics have essentially the same reading during Jun 2025 (see attached plots). And to try to answer Andrew's question about the stability of the current in the magnet, as far as epics recording goes, it seems to be at the level of about 80ppm. I asked the SL to correct the run list to show the readback values for sbb for the entire run list. There was also a correction to make to date, most of the 24h of 14-Jun were recorded as 13-Jun. Shift summaries: 23-Jun-2025 D https://logbooks.jlab.org/entry/4403887 23-Jun-2025 S https://logbooks.jlab.org/entry/4404134 24-Jun-2025 O https://logbooks.jlab.org/entry/4404338 Shift BTA: 23-Jun-2025 D https://logbooks.jlab.org/entry/4404116 23-Jun-2025 S https://logbooks.jlab.org/entry/4404327 24-Jun-2025 O https://logbooks.jlab.org/entry/4404543 (this one may be wrong, it seems that the TO recorded BANU as BNA) We will have our daily RC meeting at 1pm in the conference room on the second floor of the CH and on zoom. It seems that the currently going JLUO has a lunch break until 14:00, so it should not affect our RC meet. Zoom info: Join ZoomGov Meeting https://jlab-org.zoomgov.com/j/1616220064 Meeting ID: 161 622 0064 Passcode: 314159 --- This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/4404643 -------------- next part -------------- An HTML attachment was scrubbed... URL: From covrig at jlab.org Wed Jun 25 10:40:02 2025 From: covrig at jlab.org (covrig at jlab.org) Date: Wed, 25 Jun 2025 10:40:02 -0400 (EDT) Subject: [Halla_running] [New Logentry] RC Report for 25-Jun-2025 Message-ID: <20250625144002.237E6803BE8@elog9.acc.jlab.org> Logentry Text: -- Key updates for the past 24h (Tue Day+Swing and Wed Owl): - 14.47h actual beam delivery on target. The beam was down for 9.5h with major culprits 1L18 breaker MDP tripped off, 5.18h ( https://ace.jlab.org/dtm/incidents/14999), cooling water chillers 2.5h (https://logbooks.jlab.org/entry/4404500) - During late Swing and early Owl we did a high beam current test for about 4h, increasing the beam current in steps of 5 ?A to 35 ?A (where the daq dt was about 50%). The tests were supervised by Nilanga (in the CH). We expect the analysis to get back to us within a week as to where we stand with respect to increasing the beam current as high as acceptable. - On our side we were down for 3.1h during beam delivery, which accounts for 21.5% of wasted beam time. - Beam studies today 7-9am, then RF recovery until mid afternoon with beam recovery after. The Hall will go to CA with access to the Hall scheduled after 9am. - IHWP changed just before beam studies began this morning - Thu Day there are plans for beam studies during beam delivery to test two settings of the Wien filter (for MOLLER). The Halls are "required" to take data during this time, maybe tag it, and report about the effects (if any). MCC will announce when the Wien tests will start and stop, to give Halls time to prepare for daq. - DAQ on-call until Jul 10: primary C. H. Leung (217-693-1360), secondary Chandan (757-342-4035). - Total Accumulated Charge at Kin3 so far: 32.241 C (Goal: 100 C) - Total charge delivered on target since Jun 18 noon: 9.6 C - Total Accumulated Charge at Kin3 since June 18 noon: 7.241 C - Charge acquisition efficiency 75% Run plan: - Beam studies 7-9am, to test the MOLLER ffb, beam to HA expected, Target in Out-of-Beam position - CA for the rest of Day shift: several groups want to access the Hall (ECal, GEMs, daq), Safety Walkthrough, Moller polarimeter daq, ... - Beam recovery during Swing - Production on LH2 at 22-25 ?A on the 30 cm LH2 target cell (loop 1), updates on subsystems on: https://sbs.jlab.org/wiki/index.php/RC_meeting_and_Run_Plan,_Wednesday,_June_25th#Current_Issues Current issues: - DAQ crashes: the last 24h we wasted 3.1h (evenly between Swing and Owl, there was no beam during Day) due to them or 21.5% of actual beam delivery. - Online zero suppression performance (common-mode min values for online zero suppression were changed to zero for the first iteration of averaging until we are ready to deploy ROI readout and histogramming firmware). - HCAL calibration/trigger efficiency, almost complete, still waiting on replays to finish. Shift summaries: 24-Jun-2025 D https://logbooks.jlab.org/entry/4404564 24-Jun-2025 S https://logbooks.jlab.org/entry/4404808 25-Jun-2025 O https://logbooks.jlab.org/entry/4405030 Shift BTA: 24-Jun-2025 D https://logbooks.jlab.org/entry/4404786 24-Jun-2025 S https://logbooks.jlab.org/entry/4405035 25-Jun-2025 O https://logbooks.jlab.org/entry/4405302 We will have our daily RC meeting at 1pm in the conference room on the second floor of the CH and on zoom. It seems that the currently going JLUO has a lunch break until 14:00, so it should not affect our RC meet. Zoom info: Join ZoomGov Meeting https://jlab-org.zoomgov.com/j/1616220064 Meeting ID: 161 622 0064 Passcode: 314159 --- This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/4405404 -------------- next part -------------- An HTML attachment was scrubbed... URL: